# Subtree Mode and Applications

### v1

@@ 2 Background
Range Mode is by now a classic problem in data structures theory [KMS05, Gre+10, Cha+14, Dur+15, Dur+16, El-+18, VX20, SX20, Gu+21]. It is defined as follows.
@@
SM can be seen as a specialization of RM on leaf-colored trees. The fundamental difference is that in RM we have $\Theta(N^{2})$ distinct queries (one per interval $[i,j]$), while in SM we have $\mathcal{O}(N)$ possible query intervals that form a laminar family, that is, for every two intervals, either the intervals are disjoint or one is contained in the other. Although our $\mathcal{O}(N)$-time construction algorithm (Theorem 1) precomputes the answer of each possible query, we opted to define SM as a data structure (instead of an algorithmic) problem to be consistent with RM. Moreover, it might be possible to have a data structure of $o(N)$ size supporting (near-)optimal SM queries or tree updates. Natural variations of SM (similar to those of RM) output both the mode $c_{v}^{\max}$ and its frequency, or the least frequent color, $c_{v}^{\min}$, known as anti-mode, and its frequency $f_{v}^{\min}$. For simplicity, our algorithm is presented for the SM variation that returns the mode and its frequency but, as we show, it can be modified to compute instead the anti-mode and its frequency.
@@
A compacted trie is a trie in which each maximal branchless path is replaced with a single edge whose label is a string equal to the concatenation of that path’s edge labels. The dissolved nodes are called implicit while the preserved nodes are called explicit. A node with at least two children is called branching. For a node $v$ in a compacted trie, $\textsf{str}(v)$ is the concatenation of edge labels on the root-to-$v$ path. We define the string depth of a node $v$ as $\textsf{sd}(v)=|\textsf{str}(v)|$. The locus of a pattern $P$ is the node $v$ with the smallest string depth such that $P$ is a prefix of $\textsf{str}(v)$. The suffix tree of a string $S$, denoted by $\textsf{ST}(S)$, is the compacted trie of all suffixes of $S$ [Wei73].
@@
Lemma 1 ([Far97]).
@@
At each node of $\textsf{ST}(S)$, we can store a hash table to access an edge based on the first letter of its label. The hash tables can be constructed in $\mathcal{O}(n)$ total time with high probability and support $\mathcal{O}(1)$-time queries [Ben+24]. Spelling a pattern $P$ in a suffix tree $\textsf{ST}(S)$ then takes $\mathcal{O}(|P|)$ time [Wei73]: we start from the root, traverse down the tree edge by edge, matching as many letters as possible, until either the pattern ends, a mismatch occurs, or we land on a leaf.
@@ 3 Baselines for SM
Another way to solve SM is by observing that it can be reduced to RM by creating a single array of size $N$ that contains the color of each leaf of $\mathcal{T}$ from left to right, and then solving RM (i.e., preprocessing this array by constructing a data structure for RM on it and then querying this data structure). This can be done because each subtree of $\mathcal{T}$ in SM corresponds to a range of the array. We first recall a well-known result on RM, and then formalize this approach as follows.
Lemma 2 ([Cha+14]).
@@ 4 A Linear-Time Construction Algorithm for SM
Problems like SM, where statistics need to be computed for every subtree, are usually solved using the folklore smaller-to-larger technique (cf. [CHL07]) – also known as disjoint set union – on trees. Such an approach typically requires $\mathcal{O}(N\log N)$ time (or slightly more depending on the used data structures).
@@
This algorithm, henceforth referred to as SCM (for Splitting-Counting-Merging), outputs, for every node $v$ of $\mathcal{T}$, the mode of $v$ and its corresponding frequency.
@@
For each leaf list $\mathcal{L}_{i}$, we construct a leaf-colored tree $\mathcal{T}_{i}$ using a single color $i$. In particular, $\mathcal{T}_{i}$ is a tree with the elements of $\mathcal{L}_{i}$ as leaves and internal nodes being in one-to-one correspondence with the elements of the set $\{\textsf{LCA}_{\mathcal{T}}(\phi_{i}(u),\phi_{i}(v)):u,v\in\mathcal{L}_{i}\}$, where $\textsf{LCA}_{T}(x,y)$ denotes the lowest common ancestor of two nodes in a tree $T$. Our goal is to construct $\mathcal{T}_{i}$ in $\mathcal{O}(|\mathcal{L}_{i}|)$ time. The tree $\mathcal{T}_{i}$ can be constructed in $\mathcal{O}(|\mathcal{L}_{i}|)$ time using the algorithm of Kasai et al. [Kas+01, Section 5.2]. This algorithm simulates a traversal of any rooted tree $\mathcal{T}^{\prime}$ with no unary paths, if one has: (I) the leaf list of $\mathcal{T}^{\prime}$ (from left to right); (II) the LCA’s of adjacent leaves in the list; and (III) access to the partial order of these LCA nodes (each specified by two leaves $x,y$ such that the node is $\textsf{LCA}_{\mathcal{T^{\prime}}}(x,y)$) defined as $u<v$ if $u$ is a strict ancestor of $v$.
We have already constructed the list $\mathcal{L}_{i}$ of leaves of $\mathcal{T}_{i}$. To find the LCA’s of the leaves in $\mathcal{L}_{i}$, in a preprocessing step, we construct a data structure for answering LCA queries on $\mathcal{T}$. The construction takes $\mathcal{O}(N)$ time [BF00]. Given any two nodes $u$ and $v$ in $\mathcal{T}$, the data structure returns node $w=\textsf{LCA}_{\mathcal{T}}(u,v)$ in $\mathcal{O}(1)$ time. We scan $\mathcal{L}_{i}$, from left to right, and ask $(|\mathcal{L}_{i}|-1)$ LCA queries in $\mathcal{T}$, between the nodes associated with each pair of successive nodes of $\mathcal{L}_{i}$. This way, we compute at most $|\mathcal{L}_{i}|-1$ distinct internal nodes of $\mathcal{T}$; for each such node $w$, we create a node $v$ in $\mathcal{T}_{i}$ and set $\phi_{i}(v):=w$. Given two distinct nodes $u_{1}$ and $u_{2}$ of $\mathcal{T}_{i}$, each specified by a pair of successive leaves of which it is the LCA, we have that $u_{1}<u_{2}$ if and only if $\phi_{i}(u_{1})\neq\phi_{i}(u_{2})$ and $\textsf{LCA}_{\mathcal{T}}(\phi_{i}(u_{1}),\phi_{i}(u_{2}))=\phi_{i}(u_{1})$; these conditions can be checked in $\mathcal{O}(1)$ time. Thus, using Kasai et al.’s algorithm, we construct $\mathcal{T}_{i}$ and $\phi_{i}$ in $\mathcal{O}(|\mathcal{L}_{i}|)$ time. We have thus proved the following lemma.
@@
It seems difficult to improve on the above-mentioned $\mathcal{O}(N\log\Delta)$-time algorithm if we insist on the technique that merges two trees at a time. The crucial observation we make is that, instead of merging whole trees, we can merge the counts of the (at most) $\Delta$ nodes in $\mathcal{T}_{0},\ldots,\mathcal{T}_{\Delta-1}$ that are associated with the same node in $\mathcal{T}$ at once. Even if for a single node $v$ in $\mathcal{T}$, we have $\Delta$ nodes $u$ in $\mathcal{T}_{0},\ldots,\mathcal{T}_{\Delta-1}$ with $\phi_{i}(u)=v$, the total size of $\mathcal{T},\mathcal{T}_{0},\ldots,\mathcal{T}_{\Delta-1}$ is $\mathcal{O}(N)$, and so the running time amortizes to $\mathcal{O}(N)$.
More specifically, we make a single traversal of $\mathcal{T}$, processing nodes in a bottom-up manner. At every internal node $v$ of $\mathcal{T}$, we compute and store the maximum among the counts stored by the children of $v$ and the counts stored at all internal nodes $u$ of $\mathcal{T}_{i}$ with $\phi_{i}(u)=v$, for $i\in[0,\Delta)$ (recall that the latter counts were computed in Step 2). The base case is at the leaves, where we initialize the count to $1$ for the color of the leaf. We also store a corresponding most frequent color per node, breaking ties arbitrarily. Thus, we obtain the following:
@@
We now prove the correctness of our construction algorithm.
@@
Lemmas 4, 5, 6 and 7 imply Theorem 1. The space used by SCM is bounded by the construction time, which is $\mathcal{O}(N)$.
@@
If we want to compute the anti-mode instead of the mode for every node of $\mathcal{T}$, we can modify SCM as follows. The first two steps are identical. We then change Step 3 from a bottom-up to a top-down traversal. The base case is now at the root, where we simply take as the answer, the pair $(f,i)$, where $i$ is the anti-mode of the root of $\mathcal{T}$ and $f$ its frequency (breaking ties arbitrarily). For an arbitrary node $v$ of $\mathcal{T}$, we perform the following:
  1. (I)
    Let $(f,i)$ be the pair stored at $v$. If $\mathcal{T}_{i}$ does not have a node $u$ with $\phi_{i}(u)=v$, we push $(f,i)$ downwards to its corresponding child (i.e., the child whose subtree has $f$ leaves colored $i$) and $(0,i)$ to all other children of $v$.
  2. (II)
    If any $\mathcal{T}_{i}$, with $i\in[0,\Delta)$, has a node $u$ with $\phi_{i}(u)=v$, for each child $w$ of $u$, with count $f_{w}$, we push the pair $(f_{w},i)$ to the child of $v$ that is an ancestor of $\phi(w)$ in $\mathcal{T}$—this node can be computed in constant time after a linear-time preprocessing of $\mathcal{T}$ for level ancestor queries [BV94]—and $(0,i)$ to every other child of $v$; we ensure that $(0,\star)$ is pushed to each node at most once by storing a list of children of $v$ to which $(0,\star)$ has not been already pushed.
When all pairs are pushed downwards from $v$, each of the children of $v$ selects the pair $(f,i)$ with minimum $f$ among its list of pairs (breaking ties arbitrarily) and discards the rest. The total number of pairs pushed downwards is asymptotically linear in the total number of edges in the $\mathcal{T}_{i}$’s and $\mathcal{T}$, and thus linear in $N$. As noted, each such push can be performed in constant time and hence the running time of the algorithm is $\mathcal{O}(N)$.
@@
To see that this algorithm is correct, fix an edge $(\textsf{parent}(v),v)$ in $\mathcal{T}$ and a color $i$. We have three cases when descending from $\textsf{parent}(v)$ to $v$: either the frequency of $i$ is the same in $\textsf{parent}(v)$ and $v$; or the frequency of $i$ is non-zero in $\textsf{parent}(v)$ and zero in $v$; or the frequency of $i$ is reduced from $\textsf{parent}(v)$ to $v$ to a non-zero value. In the latter case, there exists a node $u$ in $\mathcal{T}_{i}$ with $\phi_{i}(u)=v$. Thus, if $i$ is the anti-mode of $v$, the algorithm will consider $i$ as part of its minimum computation. We thus obtain the following result.

### v2

@@ 2 Background
Range Mode is by now a classic problem in data structures theory [DBLP:journals/njc/KrizancMS05, DBLP:conf/icalp/GreveJLT10, DBLP:journals/corr/abs-1101-4068, DBLP:journals/tcs/DurocherEMT15, DBLP:journals/algorithmica/DurocherSST16, DBLP:conf/esa/El-Zein0MS18, DBLP:conf/soda/WilliamsX20, DBLP:conf/icalp/SandlundX20, DBLP:conf/icalp/Gu0WX21]. It is defined as follows.
@@
SM can be seen as a specialization of RM on leaf-colored trees. The fundamental difference is that in RM we have $\Theta(N^{2})$ distinct queries (one per interval $[i,j]$), while in SM we have $\mathcal{O}(N)$ possible query intervals that form a laminar family, that is, for every two intervals, either the intervals are disjoint or one is contained in the other. Although our $\mathcal{O}(N)$-time construction algorithm (Theorem 1) precomputes the answer of each possible query, we opted to define SM as a data structure (instead of an algorithmic) problem to be consistent with RM. Moreover, it might be possible to have a data structure of $o(N)$ size supporting (near-)optimal SM queries or tree updates. Natural variations of SM (similar to those of RM) output both the mode $c_{v}^{\max}$ and its frequency, or the least frequent color, $c_{v}^{\min}$, known as anti-mode, and its frequency $f_{v}^{\min}$; note that for a node $v$, the anti-mode $c_{v}^{\min}$ may have $f_{v}^{\min}=0$.
For simplicity, our algorithm is presented for the SM variation that returns the mode and its frequency but, as we show, it can be modified to compute instead the anti-mode and its frequency.
@@
A compacted trie is a trie in which each maximal branchless path is replaced with a single edge whose label is a string equal to the concatenation of that path’s edge labels. The dissolved nodes are called implicit while the preserved nodes are called explicit. A node with at least two children is called branching. For a node $v$ in a compacted trie, $\textsf{str}(v)$ is the concatenation of edge labels on the root-to-$v$ path. We define the string depth of a node $v$ as $\textsf{sd}(v)=|\textsf{str}(v)|$. The locus of a pattern $P$ is the node $v$ with the smallest string depth such that $P$ is a prefix of $\textsf{str}(v)$. The suffix tree of a string $S$, denoted by $\textsf{ST}(S)$, is the compacted trie of all suffixes of $S$ [DBLP:conf/focs/Weiner73].
@@
Lemma 1 ([DBLP:conf/focs/Farach97]).
@@
At each node of $\textsf{ST}(S)$, we can store a hash table to access an edge based on the first letter of its label. The hash tables can be constructed in $\mathcal{O}(n)$ total time with high probability and support $\mathcal{O}(1)$-time queries [DBLP:conf/sosa/BenderFKK24]. The total size of the hash tables is also bounded by $\mathcal{O}(n)$ [DBLP:conf/sosa/BenderFKK24]. Spelling a pattern $P$ in a suffix tree $\textsf{ST}(S)$ then takes $\mathcal{O}(|P|)$ time [DBLP:conf/focs/Weiner73]: we start from the root, traverse down the tree edge by edge, matching as many letters as possible, until either the pattern ends, a mismatch occurs, or we land on a leaf.
@@ 3 Baselines for SM
Another way to solve SM is by observing that it can be reduced to RM by defining a single array that stores the color of each leaf of $\mathcal{T}$ from left to right, and then solving RM on this array (i.e., preprocessing this array by constructing a data structure for RM on it and then querying this data structure). This can be done because each subtree of $\mathcal{T}$ in SM corresponds to a range of the array; and each range can be processed by a single query in the data structure for RM. The drawback of this approach, however, is that constructing the data structure and answering each of the $N$ queries takes $\mathcal{O}(N\sqrt{N})$ time, as shown below, whereas SM can be solved in the optimal $\mathcal{O}(N)$ time by using our algorithm. We first recall a well-known result on RM, and then formalize this approach as follows.
Lemma 2 ([DBLP:journals/corr/abs-1101-4068]).
@@ 4 A Linear-Time Construction Algorithm for SM
Problems like SM, where statistics need to be computed for every subtree, are usually solved using the folklore smaller-to-larger technique (cf. [DBLP:books/daglib/0020103]) – also known as disjoint set union – on trees. Such an approach typically requires $\mathcal{O}(N\log N)$ time (or slightly more depending on the used data structures).
@@
This algorithm, henceforth referred to as SCM (for Splitting-Counting-Merging), outputs, for every node $v$ of $\mathcal{T}$, the mode of $v$ and its frequency; see Algorithm 1 for the pseudocode.
Algorithm 1 $\textsf{SCM}(\mathcal{T},\Delta)$
[…]
2:  Construct single-color tree $\mathcal{T}_{i}$
3:end for
[…]
5:  for all internal nodes $u$ in $\mathcal{T}_{i}$ do
6:   $\textit{count}[u]\leftarrow\text{number of leaf descendants of }u$
7:  end for
8:end for
9:Initialize array mode with NN -1’s and array freq with NN 0’s $\triangleright$ Step 3
10:for all leaf nodes $v$ of color $c$ do
11:  $(\textit{freq}[v],\textit{mode}[v])\leftarrow(1,c)$
12:end for
13:for all internal nodes $v$ in a bottom-up traversal of $\mathcal{T}$ do
14:  $\mathcal{C}_{v}\leftarrow$ set of counts comprised of:
15:     $\textit{freq}[u]$ for each child $u$ of $v$ in $\mathcal{T}$
16:     $\textit{count}[u]$ for each $\mathcal{T}_{i}$ and each internal node $u$ in $\mathcal{T}_{i}$ s.t. $\phi_{i}(u)=v$
17:  $f^{\max}_{v}\leftarrow\max(\mathcal{C}_{v})$ and $c^{\max}_{v}\leftarrow$ a color with frequency $f^{\max}_{v}$
18:  $\textit{freq}[v]\leftarrow f^{\max}_{v}$ and $\textit{mode}[v]\leftarrow c^{\max}_{v}$
19:end for
20:return the arrays mode and freq
@@
For each leaf list $\mathcal{L}_{i}$, we construct a leaf-colored tree $\mathcal{T}_{i}$ using a single color $i$. In particular, $\mathcal{T}_{i}$ is a tree with the elements of $\mathcal{L}_{i}$ as leaves and internal nodes being in one-to-one correspondence with the elements of the set $\{\textsf{LCA}_{\mathcal{T}}(\phi_{i}(u),\phi_{i}(v)):u,v\in\mathcal{L}_{i}\}$, where $\textsf{LCA}_{T}(x,y)$ denotes the lowest common ancestor of two nodes in a tree $T$. Our goal is to construct $\mathcal{T}_{i}$ in $\mathcal{O}(|\mathcal{L}_{i}|)$ time. The tree $\mathcal{T}_{i}$ can be constructed in $\mathcal{O}(|\mathcal{L}_{i}|)$ time using the algorithm of Kasai et al. [DBLP:conf/cpm/KasaiLAAP01, Section 5.2]. This algorithm simulates a traversal of any rooted tree $\mathcal{T}^{\prime}$ with no unary paths, if one has: (I) the leaf list of $\mathcal{T}^{\prime}$ (from left to right); (II) the LCA’s of adjacent leaves in the list; and (III) access to the partial order of these LCA nodes (each specified by two leaves $x,y$ such that the node is $\textsf{LCA}_{\mathcal{T^{\prime}}}(x,y)$) defined as $u<v$ if $u$ is a strict ancestor of $v$.
We have already constructed the list $\mathcal{L}_{i}$ of leaves of $\mathcal{T}_{i}$. To find the LCA’s of the leaves in $\mathcal{L}_{i}$, in a preprocessing step, we construct a data structure for answering LCA queries on $\mathcal{T}$. The construction takes $\mathcal{O}(N)$ time [DBLP:conf/latin/BenderF00]. Given any two nodes $u$ and $v$ in $\mathcal{T}$, the data structure returns node $w=\textsf{LCA}_{\mathcal{T}}(u,v)$ in $\mathcal{O}(1)$ time. We scan $\mathcal{L}_{i}$, from left to right, and ask $(|\mathcal{L}_{i}|-1)$ LCA queries in $\mathcal{T}$, between the nodes associated with each pair of successive nodes of $\mathcal{L}_{i}$. This way, we compute at most $|\mathcal{L}_{i}|-1$ distinct internal nodes of $\mathcal{T}$; for each such node $w$, we create a node $v$ in $\mathcal{T}_{i}$ and set $\phi_{i}(v):=w$. Given two distinct nodes $u_{1}$ and $u_{2}$ of $\mathcal{T}_{i}$, each specified by a pair of successive leaves of which it is the LCA, we have that $u_{1}<u_{2}$ if and only if $\phi_{i}(u_{1})\neq\phi_{i}(u_{2})$ and $\textsf{LCA}_{\mathcal{T}}(\phi_{i}(u_{1}),\phi_{i}(u_{2}))=\phi_{i}(u_{1})$; these conditions can be checked in $\mathcal{O}(1)$ time. Thus, using Kasai et al.’s algorithm, we construct $\mathcal{T}_{i}$ and $\phi_{i}$ in $\mathcal{O}(|\mathcal{L}_{i}|)$ time. We have thus proved the following lemma.
@@
It seems difficult to improve on the above-mentioned $\mathcal{O}(N\log\Delta)$-time algorithm if we insist on the technique that merges two trees at a time. The crucial observation we make is that, instead of merging whole trees, we can merge the counts of the (at most) $\Delta$ nodes in $\mathcal{T}_{0},\ldots,\mathcal{T}_{\Delta-1}$ that are associated with the same node in $\mathcal{T}$ at once. More specifically, we make a single traversal of $\mathcal{T}$, processing nodes in a bottom-up manner. At every internal node $v$ of $\mathcal{T}$, we compute and store the maximum among the counts stored by the children of $v$ and the counts stored at all internal nodes $u$ of $\mathcal{T}_{i}$ with $\phi_{i}(u)=v$, for $i\in[0,\Delta)$ (recall that the latter counts were computed in Step 2). The base case is at the leaves, where we initialize the count to 11 for the color of the leaf. We also store the corresponding most frequent color per node, breaking ties arbitrarily.
Even if for a single node $v$ in $\mathcal{T}$, we have $\Delta$ nodes $u$ in $\mathcal{T}_{0},\ldots,\mathcal{T}_{\Delta-1}$ with $\phi_{i}(u)=v$, the running time of Step 3 amortizes to $\mathcal{O}(N)$ because: (1) accessing the associated node $u\in\mathcal{T}_{i},i\in[0,\Delta),$ of a node $v=\phi_{i}(u)\in\mathcal{T}$ takes $\mathcal{O}(1)$ time (e.g., when $\phi_{i}()$ is implemented by a pointer from $u$ to $v$); (2) each node of any $\mathcal{T}_{i},i\in[0,\Delta),$ is accessed once, by the definition of $\phi_{i}()$; and (3) the total size of $\mathcal{T},\mathcal{T}_{0},\ldots,\mathcal{T}_{\Delta-1}$ is $\mathcal{O}(N)$, as explained above. Thus, we obtain the following:
@@
We prove the correctness of SCM and that it takes $\mathcal{O}(N)$ time and space.
@@
Lemma 8.
The SCM algorithm takes $\mathcal{O}(N)$ time and space.
Proof.
Each step of SCM is executed once (see also Algorithm 1). Thus, the time follows from Lemmas 4, 5 and 6. The space used by SCM is bounded by the time, which is $\mathcal{O}(N)$. In particular, Steps 1 and 2 take $\mathcal{O}(N)$ space by 2 (for the leaf lists and trees) and by [DBLP:conf/latin/BenderF00] (for the LCA data structure); and Step 3 takes $\mathcal{O}(N)$ space since the total size of $\mathcal{T},\mathcal{T}_{0},\ldots,\mathcal{T}_{\Delta-1}$ is $\mathcal{O}(N)$ and we store $\mathcal{O}(N)$ counts. ∎
Lemmas 7 and 8 imply Theorem 1.
@@
In a preprocessing step, we compute, for every node $v$ of $\mathcal{T}$, the number $c(v)$ of distinct colors in the leaves of the subtree rooted at $v$, in $\mathcal{O}(N)$ time using a well-known algorithm [DBLP:conf/cpm/Hui92]. Then, using a bottom-up traversal on $\mathcal{T}$, we find the nodes $u$ of $\mathcal{T}$, such that $c(u)<\Delta$ and $c(\textsf{parent}(u))=\Delta$. We call these the border nodes of $\mathcal{T}$. Two observations are immediate. First, every border node and all its strict descendants have an anti-mode with a zero frequency. Second, the subtrees rooted at border nodes are pairwise disjoint. Using a bottom-up traversal per subtree, we compute a representative absent color per border node (and propagate it to all its descendants). Specifically, for each such subtree, we (re-)use a bit array $B$ of size $\Delta$ to record all colors encountered in its traversal and report the smallest color $i$ with $B[i]=0$. As we report one color per subtree, this takes time linear in the subtree size. (We re-set all array entries to 0 once the subtree is processed, and process the next subtree.) The total time is thus $\mathcal{O}(N)$. We then focus on the top part of $\mathcal{T}$: the set of strict ancestors of the border nodes; every such ancestor has an anti-mode with a non-zero frequency.
To this end, we modify SCM as follows. The first two steps are identical. We change Step 3 from a bottom-up to a top-down traversal. The base case is now at the root, where we simply take as the answer, the pair $(f,i)$, where $i$ is an anti-mode of the root of $\mathcal{T}$ and $f$ its frequency (breaking ties arbitrarily). For an arbitrary node $v$ of $\mathcal{T}$, we perform the following: (I) Let $(f,i)$ be the pair stored at $v$. If $\mathcal{T}_{i}$ does not have a node $u$ with $\phi_{i}(u)=v$, we push $(f,i)$ downwards to its corresponding child (i.e., the child whose subtree has $f$ leaves colored $i$). (II) If any $\mathcal{T}_{i}$, with $i\in[0,\Delta)$, has a node $u$ with $\phi_{i}(u)=v$, for each child $w$ of $u$, with count $f_{w}$, we push the pair $(f_{w},i)$ to the child of $v$ that is an ancestor of $\phi(w)$ in $\mathcal{T}$—this node can be computed in constant time after a linear-time preprocessing of $\mathcal{T}$ for level ancestor queries [DBLP:journals/jcss/BerkmanV94]. When all pairs are pushed downwards from $v$, each child of $v$ (lying on the top part of $\mathcal{T}$) selects the pair $(f,i)$ with minimum $f$ among its list of pairs (breaking ties arbitrarily) and discards the rest. The total number of pairs pushed downwards is asymptotically linear in the total number of edges in the $\mathcal{T}_{i}$’s and $\mathcal{T}$, and thus linear in $N$. As noted, each such push can be performed in constant time and hence the total running time of the algorithm is $\mathcal{O}(N)$.
@@
To see that the algorithm is correct, fix an edge $(\textsf{parent}(v),v)$ in $\mathcal{T}$ and a color $i$. We have three cases when descending from $\textsf{parent}(v)$ to $v$: either the frequency of $i$ is non-zero in $\textsf{parent}(v)$ and zero in $v$; or the frequency of $i$ is the same in $\textsf{parent}(v)$ and $v$; or the frequency of $i$ is reduced from $\textsf{parent}(v)$ to $v$ to a non-zero value. The first case is handled by the preprocessing. The second case is handled by Step 3(I). In the third case, there exists a node $u$ in $\mathcal{T}_{i}$ with $\phi_{i}(u)=v$. This case is handled by Step 3(II). To conclude, if $i$ is the anti-mode of $v$, the algorithm will consider $i$ as part of its minimum computation. We thus obtain the following result.
@@
[…]
Figure 5: The tree $\mathcal{T}$ of Fig. 4 annotated with (anti-mode, color) pairs; those of the leaves are the same as their parents.